OpenStudy (anonymous):

small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will total $ 23,496 . The variable costs will be $ 12 per book. The publisher will sell the finished product to bookstores at a price of $ 20.25 per book. How many books must the publisher produce and sell so that the production costs will equal the money from sales? @sleepyjess

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

if anyone is viewing this please message me because my computer is slow and I really need help with math thank you!

TheSmartOne (thesmartone):

\(\sf You~ were~ given:\) The one-time fixed costs will total $ 23,496 . The variable costs will be $ 12 per book. \(\sf Make~an~equation~out~of~that\)

OpenStudy (anonymous):

$23,496+12x @TheSmartOne

TheSmartOne (thesmartone):

correct

TheSmartOne (thesmartone):

x is the number of books needed to be made.

TheSmartOne (thesmartone):

\(\sf\Large 20.25= \frac{12x + 23,496}{x}\)

TheSmartOne (thesmartone):

Solve for x.

OpenStudy (anonymous):

@TheSmartOne so I just divide the 20.25=12x+23,496/x

TheSmartOne (thesmartone):

12x/x = 12 So, you have: 20.25=12+23,496/x solve for x.